### v1

# LV-UNet: A Lightweight and Vanilla Model for Medical Image Segmentation

Abstract
Although the progress made by large models in computer vision, optimization challenges, the complexity of transformer models, computational limitations, and the requirements of practical applications call for simpler designs in model architecture for medical image segmentation, especially in mobile medical devices that require lightweight and deployable models with real-time performance. However, some of the current lightweight models exhibit poor robustness across different datasets, which hinders their broader adoption. This paper proposes a lightweight and vanilla model called LV-UNet, which effectively utilizes pre-trained MobileNetv3-Large models and introduces fusible modules. It can be trained using an improved deep training strategy and switched to deployment mode during inference, reducing both parameter count and computational load. Experiments are conducted on ISIC 2016, BUSI, CVC-ClinicDB, CVC-ColonDB, and Kvair-SEG datasets, achieving better performance compared to the state-of-the-art and classic models.
Index Terms:
Medical image segmentation, lightweight, vanilla, MobileNetv3-Large, deep training strategy
I Introduction
Medical image segmentation, which aims to accurately delineate anatomical structures or abnormalities, is essential in disease diagnosis and treatment planning. Unlike traditional methods that relied on manual or semiautomatic methods, which were time-consuming, subjective, and prone to interobserver variability, computer-assisted techniques can be more efficient. With advances in deep learning, which utilize large annotated datasets and well-designed neural networks to learn complex image representations and automatically infer pixel-level labels, medical image segmentation has been revolutionized.
Recently, medical imaging solutions have been transitioning from laboratory settings to bedside environments. Termed point-of-care imaging, this approach involves performing testing and analysis directly alongside the patient, with the aim of improving patient care. This shift expects medical AI models to be more real-time and lightweight. For example, task shifting for point-of-care ultrasound (POCUS) includes applications in obstetrics, gynecology, emergency medicine, infectious diseases, and cardiac, abdominal, and pulmonary conditions. At the primary care level in resource-limited settings, it has the potential to expand diagnostic imaging capacity and lead to meaningful health outcomes [1]. AI systems for the point-of-care require considerations for lightweight design, deployability, and real-time performance. For example, Dulmage et al. [2] developed a point-of-care and real-time AI system to support clinician diagnosis of a wide range of skin diseases. Lightweight and fast segmentation is crucial in these scenarios, as it ensures the AI system can quickly and accurately delineate areas of interest, facilitating efficient operation on resource-constrained devices and providing precise, timely results.
UNet [3] is a classic medical image segmentation model consisting of an encoder and a decoder with skip-connections, showing a significant improvement compared to previous attempts. UNet’s variants like Attention-UNet [4], U-Net++ [5], ResUnet [6], Unet 3+[7] and so on have achieved great success. Due to the impressive performance of Transformers in visual tasks in recent years, Transformer-based U-Net variants like Trans-UNet[8], Swin-UNet [9], Unetr[10] and so on have gained wider recognition and have shown remarkable effectiveness. However, these models entail significant computational overhead and a large number of parameters, making them difficult to use in point-of-care applications. Thus, more and more effort is paid to designing lightweight models in medical image segmentation scenarios. Although existing lightweight medical image segmentation models have made significant progress, they often lack robustness across different datasets. How to address these issues by designing a model with reduced computational overhead, fewer parameters, faster inference time, and maintained performance, while also ensuring robustness across various datasets still remains a challenge.
The model design of this paper is based on the following ideas: 1) Utilizing a pre-trained lightweight model in the encoder of a segmentation model is crucial, ensuring the model’s robustness across different datasets; 2) Combining pre-trained and expansion modules forms a straightforward approach to designing lightweight medical image segmentation models; 3) Introducing fusion modules enables further reduction of parameter count and computational load during the inference phase. We propose a UNet variant called LV-UNet with a lightweight and vanilla model design containing pre-trained weights from MobileNetv3-Large and fusible expansion modules, and also adopt the fusion method in VanillaNet and improve the deep training strategy. Experiments are conducted on the ISIC 2016 [11], BUSI [12], CVC-ClinicDB [13], CVC-ColonDB [14], and Kvair-SEG datasets [15], whose results show that the presented method has competitive performances compared to the state-of-the-art and classic methods.
II Related Works
II-A Lightweight medical image segmentation models
Significant efforts have been devoted to designing lightweight models for medical image segmentation, with a primary focus on variants of U-Net. Valanarasu et al. [16] proposed UNeXt, a convolutional MLP-based network for image segmentation, featuring an initial convolutional stage and a latent MLP stage with tokenized MLP blocks that project convolutional features, reducing parameters and computational complexity while improving segmentation performance by shifting input channels to learn local dependencies. Ruan et al.[17] proposed a MAL-UNet that integrates several different modules: DGA for global and local feature extraction, IEA for dataset characterization and sample connectivity enhancement, CAB for multi-stage feature fusion generating channel-axis attention maps, and SAB for spatial-axis attention map generation on multi-stage features. Ruan et al. [18] proposed EGE-UNet, integrating Group multi-axis Hadamard Product Attention (GHPA) and Group Aggregation Bridge (GAB) modules to extract diverse pathological information and fuse multi-scale features effectively, significantly reducing the model size. Liu et al. [19] proposed Rolling-UNet, combining CNN with MLP to improve medical image segmentation by effectively capturing both local features and long-distance dependencies. Li et al. [20] introduced U-KAN, an enhanced U-Net variant integrating Kolmogorov-Arnold networks (KANs) [21] to improve accuracy and interpretability in medical image segmentation and diffusion models, which achieved superior performance with reduced computational cost compared to traditional U-Net approaches.
II-B VanillaNet
”Vanilla” is a term often used metaphorically to describe something as plain, ordinary, or unadorned without anything redundant. Pursuing minimalism in design, VanillaNet [22] adopts a simple architecture that prioritizes computational efficiency and information preservation. By combining minimalistic operations, batch normalization, and activation functions, VanillaNet achieves competitive performance without excessive complexity. The deep training strategy and the series-informed activation function are keys to maintaining non-linearity in shallow networks.
II-C MobileNetv3
MobileNetv3 is a state-of-the-art convolutional neural network designed for mobile and edge devices, combining high performance with efficiency. It builds on MobileNetv1 and MobileNetv2 [23] by integrating advanced techniques such as depthwise separable convolutions, inverted residuals with linear bottlenecks, and Squeeze-and-Excitation (SE) blocks. In addition, it introduces the H-Swish activation function, which improves computational efficiency and performance. MobileNetv3 uses neural architecture search (NAS) to optimize its structure for different performance and efficiency needs, resulting in two variants: MobileNetv3-Large for higher accuracy and MobileNetv3-Small for resource-constrained environments. It also incorporates training techniques like dropout, label smoothing, and mixup for better generalization.
III Motivation
III-A The Role of Pre-training
Pre-training in large datasets allows models to learn rich and generalized feature representations, which significantly improves their performance in various tasks. These learned features capture important patterns and nuances in the data that may not be easily obtained through training on smaller datasets or from scratch. Lightweight models have fewer parameters, which can restrict their ability to learn and generalize from data effectively. Therefore, even though introducing complex changes to the architectures of model might seem beneficial, these modifications often do not provide enough improvement to outweigh the substantial performance boost gained from using pre-trained models. On the other hand, while complex model structures might offer improvements on specific datasets, such benefits are usually inconsistent. These intricate designs may overfit particular data characteristics, leading to poor generalization in different or unseen datasets. Relying on pre-trained weights is a more reliable approach to ensure the model’s overall performance and robustness.
III-B Efficient Model Deployment through Reparameterization
During the training phase, a more complex model can be utilized to ensure optimal learning and feature extraction. This model may include redundancy in parameters and use techniques like dropout and batch normalization for improved generalization. Once training is complete, the model undergoes reparameterization for the inference phase. This involves simplifying the model by compressing and fusing parameters, such as merging multiple convolutional filters into fewer ones. This reparameterized model maintains performance while significantly reducing computational and memory requirements, ensuring efficient deployment in resource-constrained environments.
IV Methods
IV-A Overall Architecture
The model architecture contains an encoder, a decoder, and final output layers, which can be seen in Figure 1. The encoder includes pre-trained blocks from MobileNetv3-Large, as well as fusible encoder modules. The decoder consists of fusible decoder modules. There are skip connections between the corresponding layers in the encoder and decoder, and to avoid significantly increasing the number of parameters, element-wise addition is used instead of concatenation. The dimensions of each stage are [16, 24, 40, 80, 160, 240, 480].
Figure 1: The architecture of LV-UNet: the basic modules are pre-trained MobileNetv3-Large blocks(the initial convolution stage and the group \@slowromancapi@ to \@slowromancapiii@ (the first inverted residual block to ninth), fusible encoder blocks, fusible decoder blocks, skip-connections and the output block
IV-B Pre-trained Modules in Encoder
LV-UNet uses MobileNetv3-Large from the initial convolutional stage to the ninth inverted residual block for the encoder block of LV-UNet. The initial convolutional stage receives the input image and performs basic feature extraction, and the others contain inverted residual blocks and depth-wise separable convolution layers, each with different convolution kernels and activation functions, progressively extracting more abstract features. We name fifteen inverted residual blocks in MobileNetv3-Large model as $r1$ to $r15$. These inverted residual blocks are grouped as follows: $r1$ and $r2$ as the first group -group \@slowromancapi@ , $r3$, $r4$, and $r5$ as the second group -group \@slowromancapii@, and $r6$, $r7$, $r8$, and $r9$ as the third group -group \@slowromancapiii@. We do not use the entire MobileNetv3-Large without fully connected layers, to allow flexible design of subsequent expansion modules. If we included inverted Residual block $r10$ through $r15$ after $r9$, the number of downsampling operations (and thus the image size constraints) would limit the fusion modules.
IV-C Fusible Modules
IV-C1 Architecture
The training mode of a fusible encoder block contains a convolutional layer using a $1\times 1$ kernel with stride 1, a batch normalization operation, a Leaky ReLU function, a convolutional layer using a $1\times 1$ kernel, a pooling layer and a designed activation layer. The resolutions of output features of an encoder block are half of the input. In the deployment mode, the block only contains a convolutional layer, a pooling layer and the designed activation layer. A fusible decoder block is similar, where the pooling operation is replaced by an upsampling operation. The resolutions of output features of a decoder block are two times of the input. The architectures of the fusible blocks in the training and deployment modes are shown in Figure 2. The method to transfer the training mode to the deployment mode during the inference stage will be introduced in the section ”Deep Training Strategy and Deployment Mode”.
(a) The architecture of the fusible encoder block in the training and deployment modes
(b) The architecture of the fusible decoder block in the training and deployment modes
Figure 2: The architecture of the fusible blocks in the training and deployment modes
IV-C2 Non-Linear Activation Layer
The Non-Linear Activation Function in LV-UNet follows the design in VanillaNet. Consider an input feature tensor, denoted as $x\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ represent the height, width, and number of channels, respectively. The activation function is formulated as follows:
$A_{s}\left(x_{h,w,c}\right)=\sum_{i,j\in\{-n,n\}}a_{i,j,c}A\left(x_{i+h,j+w,c}+b_{c}\right)$ (1)
where $A_{s}\left(x_{h,w,c}\right)$ denotes the output of the activation function at position $(h,w,c)$ in the input tensor $h$ ranges from $1$ to $H$, $w$ ranges from $1$ to $W$, and $c$ ranges from $1$ to $C$. The function aggregates inputs from neighboring positions $(i+h,j+w,c)$ within a range defined by $-n$ to $n$. The weights $a_{i,j,c}$ control the contribution of each input, and $A(\cdot)$ represents the activation function applied to each input, typically a nonlinear function like ReLU. Furthermore, $b_{c}$ represents a bias term specific to each channel $c$. The Non-Linear Activation Function is followed by a convolution and a batch normalization operation. By utilizing this formulation, LV-UNet can capture and incorporate contextual information from neighboring positions, enabling the model to learn and represent complex patterns and non-linear relationships within the input data.
IV-C3 Deep Training Strategy and Deployment Mode
VanillaNet [22] adopts a deep training strategy in the training process, and it is also applied in the LV-UNet training. At the start of the training process, two convolutional layers utilize an activation function. During the training epochs, the activation function is progressively adjusted to resemble an identity mapping. At the end of the training, the two convolutions can seamlessly combine into a single convolution, resulting in reduced inference time. We use a Leaky ReLU function to control it:
[…]
$a=1-cos(\frac{\pi e}{2E})$ (3)
where $e$ is the current epoch and $E$ is the number of deep training epochs. At the beginning of the training, the activation function is a ReLU function, while at the end of the training, it is an identity mapping function. The nonlinearity gradually decreases. Unlike the design in [22] that defines $a$ as $a=\frac{e}{E}$, the designed method can make the transition from linear to non-linear show a trend of starting slow and accelerating, which aids in the convergence of training.
Following the method used in VanillaNet’s training, the weight and bias matrices of a single convolution merged from the batch normalization layer and their preceding convolution in fusible blocks are:
$W_{i}^{\prime}=\frac{\gamma_{i}}{\sigma_{i}}W_{i},B_{i}^{\prime}=\frac{\left(B_{i}-\mu_{i}\right)\gamma_{i}}{\sigma_{i}}+\beta_{i}$ (4)
where $W\in R^{C_{\text{out }}\times\left(C_{\text{in }}\times k\times k\right)}$, $B\in R^{\text{out }}$ are the weight and bias matrices of the convolutional kernel, $\gamma,\beta,\mu,\sigma\in R^{\text{out }}$, are the scale, shift, mean and variance in batch normalization, subscript $i\in\left\{1,2,\ldots,C_{out}\right\}$ denotes the value in $i$-th output channels.
Then the two 1$\times$1 convolution layers in the fusible blocks can be merged as a single convolution layer:
$y=W^{1}*\left(W^{2}*x\right)=W^{1}\otimes W^{2}\otimes\operatorname{im}2\operatorname{col}(x)$ (5)
$=\left(W^{1}\otimes W^{2}\right)*X$
where * means the convolution operation, $\otimes$ represents the matrix multiplication, $W^{1}$ and $W^{2}$ are the weight matrix of two convolution layers, and $X\in$ $R^{\left(C_{in}\times 1\times 1\right)\times\left(H^{\prime}\times W^{\prime}\right)}$ is derived from the im2col operation to transform the input into a matrix corresponding to the shape of the kernel.
By using the deep training strategy, LV-UNet can be switched to deployment mode in the inference stage, which reduces parameters and computational requirements a lot.
V Experiments
V-A Datasets
TABLE I: Comparison Experimental Results on ISIC 2016, BUSI, CVC-ClinicDB, CVC-ColonDB and Kvasir-SEG datasets
Methods ISIC 2016 BUSI CVC-ClinicDB CVC-ColonDB Kvasir-SEG
IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
UNet 0.8401 0.9116 0.6218 0.7575 0.7649 0.8638 0.7425 0.8387 0.7327 0.8440
Attention-UNet 0.8430 0.9137 0.6445 0.7755 0.7955 0.8839 0.7523 0.8487 0.7449 0.8512
UNet++ 0.8375 0.9097 0.6211 0.7568 0.7951 0.8832 0.7246 0.8294 0.7565 0.8598
UNext 0.8505 0.9181 0.6122 0.7437 0.7709 0.8660 0.6613 0.7774 0.6880 0.8115
MALUNet 0.8374 0.9098 0.5875 0.7318 0.4994 0.6565 0.5361 0.6860 0.5724 0.7219
EGEUNet 0.8380 0.9103 0.5571 0.6949 0.5190 0.6774 0.5744 0.7098 0.5342 0.6928
Rolling-UNet-M 0.8472 0.9160 0.6132 0.7434 0.7342 0.8435 0.5382 0.6773 0.6990 0.8211
U-KAN 0.8517 0.9192 0.6368 0.7668 0.8243 0.9017 0.7367 0.8353 0.7338 0.8438
MobileNetv3-Large-UNet 0.8602 0.9240 0.6226 0.7504 0.8347 0.9087 0.7802 0.8700 0.7804 0.8747
LV-UNet 0.8595 0.9235 0.6489 0.7785 0.8514 0.9183 0.7964 0.8808 0.8069 0.8917
LV-UNet (deployment) 0.8593 0.9234 0.6347 0.7642 0.8272 0.9045 0.7975 0.8811 0.7956 0.8840
V-A1 ISIC2016
ISIC 2016 [11] is a dataset of dermoscopic images of skin lesions for Skin Lesion Analysis toward Melanoma Detection Challenge at the International Symposium on Biomedical Imaging (ISBI) 2016. In this paper, we used the lesion segmentation dataset, which contains original images and binary lesion masks. The training set has 900 pairs of original images and corresponding masks. In the testing set, there are 379 pairs of images and masks. The training set is split into a new training set and a validation set with the random state 42.
V-A2 BUSI
Breast Ultrasound Images Dataset (Dataset BUSI) [12] is a dataset of breast ultrasound images collected from women between 25 and 75 years old. The dataset consists of 780 pairs of images in normal, benign, and malignant, and corresponding masks. Unlike the setting in [16], we include all three classes in the dataset, and the test set is split with a ratio of 0.2 with a random state of 42. Then the rest of the dataset is split into the training set and the validation set with a ratio of 0.2 with a random state of 42.
V-A3 CVC-ClinicDB
CVC-ClinicDB [13] is a database of frames extracted from colonoscopy videos that contain several examples of polyps. The dataset consists of 612 pairs of images and corresponding masks. The test set is split with a ratio of 0.2 with a random state of 42. Then the rest of the dataset is split into the training set and the validation set with a ratio of 0.2 with a random state of 42.
V-A4 CVC-ColonDB
The CVC-ColonDB dataset [14] includes 380 images, each with a resolution of 574$\times$500 pixels. These images were selected from a pool of 12,000 images obtained from 15 short colonoscopy videos, with only these 380 images being annotated. The test set is split with a ratio of 0.2 with a random state of 42. Then the rest of the dataset is split into the training set and the validation set with a ratio of 0.2 with a random state of 42.
V-A5 Kvasir-SEG
The Kvasir-SEG dataset [15] comprises 1000 polyp images and their corresponding ground truth masks with resolutions ranging from 332$\times$487 to 1920$\times$1072 pixels. The test set is split with a ratio of 0.2 with a random state of 42. Then the rest of the dataset is split into the training set and the validation set with a ratio of 0.2 with a random state of 42.
Figure 3: Example segmentation results of different models on ISIC 2016, BUSI, CVC-ClinicDB, CVC-ColonDB and Kvasir-SEG.
V-B Implementation Details
The experiments are all performed on PG500-216(V-100) with 32 GB memory. The total training epochs are 300, and the batch size in training, validation, and testing is 8. The resolution of the input images is resized to 256$\times$256. The optimizer used is ADAM [24]. The initial learning rate is 0.001, and a CosineAnnealingLR [25] scheduler is used. The minimum learning rate is 0.00001. Only rotation by 90 degrees clockwise for random times, random flipping, and normalization methods are used for data augmentation. The evaluation metrics in validation are $IOU$ of the lesions. The loss function used is a mixed loss combining binary cross entropy (BCE) loss and dice loss [26]:
$\mathcal{L}=0.5BCE(\hat{y},y)+Dice(\hat{y},y)$
We also train, validate and test UNet, Attention UNet, UNet++, UNext, MLUNet, EGE-UNet, Rolling-UNet-M, U-KAN and MobileNetv3-UNet (UNet with MobileNetv3 as the backbone) for comparison. UNet uses double convolution layers in each stage. The dimensions of each stage in UNet, Attention UNet and UNet++ are [32, 64, 128, 256, 512], while the UNext, MLUNet, EGE-UNet, Rolling-UNet-M, and U-KAN use the original design in their papers. MobileNetv3-UNet adds the output of each encoder stage to the blocks in the decoder.
For other experiments, we select the model that performs the best on the validation set for inference on the test set. However, the Leaky ReLU only transitions to an identity mapping function at the end of the training. Therefore, we choose the model from the last epoch for inference on the test set, for the LV-UNet using the deep training strategy.
V-C Results
The comparison experimental results can be seen in Table \@slowromancapi@, from which we can see LV-UNet shows competitive performances compared to the baselines. The deployment mode of LV-UNet can still achieve good results while reducing parameters a lot. In Table \@slowromancapii@, the number of parameters and GFLOPs of different models are compared, which shows that the model sizes and computation complexities of LV-UNet and its deployment mode are suitable for edge device and point-of-care scenarios. And some example segmentation results can be seen in Figure 3.
TABLE II: Parameters and GFLOPs of Different Models
Methods Parameters GFLOPs
UNet 7.8M 12.15
Attention-UNet 8.7M 16.79
UNet++ 9.2M 34.77
UNext 1.5M 0.58
MALUNet 0.2M 0.09
EGEUNet 53.4K 0.08
Rolling-UNet-M 7.1M 0.83
U-KAN 25.4M 10.5
MobileNetv3-Large-UNet 2.0M 0.51
LV-UNet 0.9M 0.22
LV-UNet (deployment) 0.5M 0.20
TABLE III: Experimental Results of Different Numebers of pre-trained Layers
Methods BUSI CVC-ClinicDB Kvasir-SEG
IoU Dice IoU Dice IoU Dice
Combination 1 0.6308 0.7548 0.8531 0.9197 0.8063 0.8901
Combination 2 0.6489 0.7785 0.8514 0.9183 0.8069 0.8917
Combination 3 0.6332 0.7625 0.8324 0.9080 0.7971 0.8856
TABLE IV: Parameters and GFOPs of Different Numbers of pre-trained Layers
      Methods        Parameters       GFLOPs
      Combination 1       2.8M       0.34
      Combination 2       0.9M       0.22
      Combination 3       0.8M       0.17
V-D Ablation Studies
The Number of Pre-trained Layers We explore the number of pre-trained layers of MobileNetv3-Large used in LV-UNet, and propose three possible combinations in the encoder:
  • $\bullet$
    Combination 1: From the initial convolutional stage to the inverted residual block $r14$ in MobileNetv3-Large and two encoder fusible blocks
  • $\bullet$
    Combination 2: From the initial convolutional stage to the inverted residual block $r9$ in MobileNetv3-Large and three encoder fusible blocks
  • $\bullet$
    Combination 3: From the initial convolutional stage to the iinverted residual block $r15$ in MobileNetv3-Large and four encoder fusible blocks
TABLE V: Experimental Results of Different Numebers of Series Informed Activation Functions
Methods BUSI CVC-ClinicDB Kvasir-SEG
IoU Dice IoU Dice IoU Dice
0 0.6147 0.7506 0.8013 0.8883 0.7891 0.8803
1 0.6489 0.7785 0.8514 0.9183 0.8069 0.8917
2 0.6402 0.7731 0.8369 0.9098 0.7695 0.8649
3 0.6370 0.7611 0.8541 0.9198 0.8015 0.8847
4 0.6423 0.6423 0.8495 0.9178 0.7755 0.8707
The comparison results can be seen in Table \@slowromancapiii@, and the parameters and GFLOPs can be seen in Table \@slowromancapiv@. The results show that the LV-UNet setting ( combination 2) can achieve better performance on BUSI and Kvasir-SEG datasets and achieve a very close result on CVC-ClinicDB dataset. Meanwhile, compared to combination 1, combination 2 reduces the model size and computational complexities significantly.
The Number of Series-Informed Activation Functions In experiments in [22], the increase of the number of Series-Informed Activation Functions can lead to improvements in VanillaNet’s classification accuracy. However, in our experiments, no significant relationship was observed between the number of series-informed activation functions and the segmentation accuracy of LV-UNet. But the results can suggest that using an appropriate number of Series-Informed Activation Functions may potentially enhance segmentation accuracy. The comparison results can be seen in Table \@slowromancapv@.
TABLE VI: Comparison of adding and concatenation methods for Skip-Connection
Methods BUSI CVC-ClinicDB Kvasir-SEG
IoU Dice IoU Dice IoU Dice
Concatenation 0.6088 0.7400 0.8392 0.9119 0.8047 0.8906
Adding 0.6489 0.7785 0.8514 0.9183 0.8069 0.8917
TABLE VII: Comparison Experimental Results of Different Nonlinearity Decreasing Methods for LV-UNet (Delopyment Mode)
Methods ISIC 2016 BUSI CVC-ClinicDB CVC-ColonDB Kvasir-SEG
IoU Dice IoU Dice IoU Dice IoU Dice IoU Dice
$a={e}/{E}$ 0.8575 0.9223 0.6293 0.7609 0.8300 0.9053 0.7983 0.8810 0.7952 0.8834
$a=1-cos({\pi e}/{2E})$ 0.8593 0.9234 0.6347 0.7642 0.8272 0.9045 0.7975 0.8811 0.7956 0.8840
The Way of Skip-Connections We evaluated the two kinds of Skip-Connections methods: adding and concatenation. The adding operation in Skip-Connections outperforms concatenation on all three datasets while reducing parameters and computational complexities. The comparison results can be seen in Table \@slowromancapvi@.
Deep Training Strategies We also compared diffrent nonlinearity decreasing methods: $a=1-cos({\pi e}/{2E})$ and the design in [22] $a={e}/{E}$. The comparison results Table \@slowromancapvii@ show that performances are very close and our method achieves better results on more datasets.
VI Conclusions
This paper proposes a lightweight and vanilla model named LV-UNet, which utilizes a pre-trained MobileNetv3-Large model and introduces fusible modules. The model can be trained using an improved deep training strategy and switched to deployment mode during inference, reducing both parameter count and computational load. Experiments are conducted on ISIC 2016, BUSI, CVC-ClinicDB, CVC-ColonDB and Kvair-SEG datasets, achieving better performance compared to state-of-the-art and classic models. And the effect of the number of pre-trained layers, the number of Series-Informed Activation Functions, the way of skip-connections and deep training strategies are carefully studied. The research presents a general way to design a lightweight model for medical image segmentation: the combination of pre-trained models and fusible modules, which may provide some hints for further studies.
Acknowledgement
This work was supported by a Grant from The National Natural Science Foundation of China (No. 62103363).